A. Prepare Me-me Young Company's Common-size Statement of Financial Performance analysis for the year ended, 2020 given the following information:



-Cost of Goods Sold - 460,000.00



-Gross Profit - 640,000.00



-Net Income - 260,000.00



-Net Sales - 1,100,000.00



-Operating Expenses - 375,000.00



-Operating Income - 265,000.00



-Other Expenses - 5,000.00

1
Expert's answer
2021-11-18T14:44:01-0500

Question 266377

Me-me Young Company's Common-size Statement of Financial Performance analysis for the year ended 2020


Net sales = "1100000\/1100000*100=100%"%

Cost of Goods Sold ="460000\/1100000*100=41.82%"%

Gross profit = "640000\/1100000*100=58.18%"%

Operating expenses = "375000\/1100000*100=34.09%"%

Other expenses ="5000\/1100000*100=0.45%"%

Operating income = "265000\/1100000*100=24.09%"%

Net income = "260000\/1100000*100=23.64%"%